# Reward Inside the Model: A Lightweight Hidden-State Reward Model for LLM’s Best-of-N sampling

### v1

@@ 4.6 Logit-Based Training for Closed-Source LLMs
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
@@
While ELHSR offers significant advantages in terms of efficiency and performance, making it well-suited for deployment by large model providers and end-users, a limitation arises when dealing with closed-source LLMs where hidden states are inaccessible. To overcome this constraint, we leverage the fact that some closed-source models still provide access to their logits, such as GPT-3.5-turbo (OpenAI, 2023a) and GPT-4 (Achiam et al., 2023), as illustrated in Finlayson et al. (2024) and OpenAI (2023b). We train ELHSR directly on these logits, resulting in a model with a parameter count that is orders of magnitude smaller than traditional reward models - and even smaller than ELHSR trained on all layers. This allows for effortless deployment on personal devices such as smartphones and personal computers. As demonstrated in Table 4 and Appendix Table 9, the resulting logit-based ELHSR retains strong performance. Furthermore, as shown in Table 1, its performance even surpasses that of many baselines. These results highlight the versatility and adaptability of ELHSR, making it a viable solution even when access to hidden states is restricted.
4.7 Combining ELHSR with External Reward Models
@@
Figure 5: Average performance of combined ELHSR and external reward models across different datasets and task-performing models, for both rank selection and scaled averaging. See Appendix Figure 8 for full details.
@@
ELHSR offers a highly parameter-efficient approach to reward modeling, enabling its seamless integration with existing external reward models. Given ELHSR’s minimal parameter footprint (less than 0.005% of baseline models), the additional computational overhead introduced by incorporating ELHSR is negligible compared to relying solely on the external reward model. We explore two straightforward methods for combining ELHSR with external rewards: rank selection and scaled averaging. In rank selection, we choose the sample with the lowest (i.e. best) rank as determined by either ELHSR or the external reward model. In scaled averaging, we average the rewards from both models after normalizing them to the [0, 1] range. Detailed descriptions of these methods are provided in Appendix B.
@@
We empirically evaluated the effectiveness of these combination strategies by integrating ELHSR with various traditional reward models across different datasets and task-performing models, assessing performance using best-of-N sampling. Figure 5 presents the average results across different datasets for both ranking and scaling methods, applied to various task-performing models. Detailed results for each dataset and task-performing model can be found in Appendix Figure 8. Our findings indicate that combining ELHSR with external reward models can lead to further improvements in accuracy, particularly on the GSM8K and AQuA-RAT datasets. The MATH dataset did not exhibit significant performance gains, potentially because ELHSR already significantly outperforms the external reward models on this benchmark. This performance boost is likely due to ELHSR and external reward models capturing different types of errors, resulting in improved overall accuracy-particularly with a higher number of best-of-N reasoning paths. This combination strategy offers a pathway to achieve further performance gains without incurring significant computational overhead.
4.8 Analysis on ELHSR
@@
To investigate the mechanism of ELHSR, we present illustrative examples of its scoring of reasoning paths in Appendix F. The examples with the highest and lowest gating $\times$ reward values are highlighted, along with their corresponding numerical scores. Notably, ELHSR assigns relatively high or low gating $\times$ reward values to specific numerical and mathematical symbols, as well as particular tokens such as “answer”, “boxed” and the special token ‘<|eot_id|>’. This behavior suggests that ELHSR possesses the ability to identify and weigh key components within the reasoning process.
5 Related Work
Reward Modeling for Language Models’ Reasoning
The rapid advancement of Large Language Models (LLMs) in complex reasoning tasks (Wei et al., 2022; Chen et al., 2025; Hao et al., 2023; Kumar et al., 2025; Xia et al., 2025; Hao et al., 2024a) has highlighted the pivotal role of reward modeling in both training and post-training inference (Cui et al., 2024; Cao et al., 2024; Lambert et al., 2025; Sun et al., 2024; Scheurer et al., 2023; Ouyang et al., 2022). However, reward modeling faces several challenges. These include high computational costs (Guo et al., 2025; Namgoong et al., 2024; Zhang et al., 2024) and potentially suboptimal performance (Wen et al., 2025; Gao et al., 2023; Coste et al., 2024). Although prior work has proposed various approaches to mitigate these issues (Chan et al., 2024; Fisch et al., 2024; Meng et al., 2024), achieving a satisfactory balance between efficiency and performance remains a challenge. This work introduces a novel perspective by leveraging internal representations to achieve a significant improvement in efficiency with satisfactory performance.
The rapid development of LLMs has spurred considerable research into their internal representations (Ferrando et al., 2024; Li et al., 2023a; Gur-Arieh et al., 2025; Pan et al., 2024; Gupta et al., 2024; Nanda et al., 2023). Hidden states of LLMs have been effectively utilized in diverse fields, including safety alignment (Zou et al., 2023; Barez et al., 2025; Zhou et al., 2024; Jiang et al., 2025; Wichers et al., 2024), faithfulness and hallucination detection (Burns et al., 2023; Li et al., 2023b; Qi et al., 2024; Ferrando et al., 2025), knowledge editing (Meng et al., 2022; Yu et al., 2024; Fang et al., 2025; Gu et al., 2024b), knowledge distillation or transfer (Dasgupta and Cohn, 2025; Wu et al., 2024) and reasoning within the latent space (Wang et al., 2025; Hao et al., 2024b). By contrast, we effectively leverage the hidden states in intermediate layers of LLMs as a direct signal for reward modeling in reasoning tasks.
6 Conclusion
@@
In this paper, we presented the Efficient Linear Hidden State Reward (ELHSR) model, a novel reward modeling approach for large language models (LLMs) that leverages the rich information in the LLM’s internal hidden states with significant efficiency and efficacy, and can be scaled up efficiently in the training and inference process. We also demonstrated that comparable performance can be achieved by training ELHSR on a subset of layers or even on logits alone, further lowering computational costs and making it more versatile. The success of employing ELHSR in combination with standard reward models further illustrates its versatility and potential.
The linear nature of ELHSR opens up exciting avenues for future research. Its simplicity may contribute to improved interpretability of LLMs by providing insights into their reasoning process. Furthermore, the ELHSR framework could potentially be extended to multi-modal LLMs, enabling the development of efficient reward models for a broader range of applications. Additionally, future work could explore the use of ELHSR to directly guide the reinforcement learning training of LLMs. We believe that ELHSR represents a significant step towards more efficient and effective reward modeling for LLMs, paving the way for future advancements in the field.
Acknowledgments and Disclosure of Funding
@@
We gratefully acknowledge Mrs. Jiahe Jin from Zhiyuan College, Shanghai Jiao Tong University, for her invaluable assistance in creating the initial version of our illustrations. We also thank Mr. Zhihui Xie from The University of Hong Kong for offering precious suggestions.
@@
[…]
Appendix A Baseline Reward Models Description
@@
  • EurusRM-7B: https://huggingface.co/openbmb/Eurus-RM-7b. EurusRM-7B is a 7B parameter reward model trained on a mixture of UltraInteract [Yuan et al., 2025], UltraFeedback [Cui et al., 2024], and UltraSafety datasets [Guo et al., 2024], with a focus on improving reasoning performance. EurusRM-7B stands out as the best 7B RM overall and achieves similar or better performance than much larger baselines. Particularly, it outperforms GPT-4 in certain tasks.
  • Skywork-Reward-Llama-3.1-8B-v0.2: https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B-v0.2. This 8B parameter reward model is built on the Llama-3 architecture [Grattafiori et al., 2024] and trained on a curated dataset of 80K high-quality preference pairs. It excels at handling preferences in complex scenarios, including challenging preference pairs, and span various domains such as mathematics, coding, and safety. As of October 2024, Skywork-Reward-Llama-3.1-8B-v0.2 ranks first among 8B models.
  • Starling-RM-7B: https://huggingface.co/berkeley-nest/Starling-RM-7B-alpha. Starling-RM-7B is a reward model trained from Llama2-7B-Chat [Touvron et al., 2023], following the method of training reward models in the instructGPT paper [Ouyang et al., 2022]. It is trained with the Nectar preference dataset [Zhu et al., 2024], which is based on GPT-4’s preferences.
  • UltraRM-13B: https://huggingface.co/openbmb/UltraRM-13b. UltraRM is a reward model initialized by Llama2-13B [Touvron et al., 2023] and fine-tuned on the UltraFeedback dataset [Cui et al., 2024]. It achieves state-of-the-art performance among open-source reward models on public preference test sets.
Appendix B Detailed Approach for Combining ELHSR with External Reward Models
In the Best-of-N (BoN) setting, we explore two strategies for combining the ELHSR reward ($R_{ELHSR}$) with an external reward signal ($R_{ext}$), typically from a traditional reward model. We have $N$ reasoning paths.
  1. 1.
    Ranking-based Combination: Let $rank_{ELHSR}(i)$ and $rank_{ext}(i)$ be the ranks of the $i$-th reasoning path according to ELHSR and the external reward model, respectively (lower rank is better). The relative rank is:
    $rank_{rel}(i)=\max(rank_{ELHSR}(i),rank_{ext}(i))$ (6)
@@ Appendix B Detailed Approach for Combining ELHSR with External Reward Models
    $\tilde{R}_{ELHSR}(i)=\frac{R_{ELHSR}(i)-\min_{j}R_{ELHSR}(j)}{\max_{j}R_{ELHSR}(j)-\min_{j}R_{ELHSR}(j)}$ (8)
    $\tilde{R}_{ext}(i)=\frac{R_{ext}(i)-\min_{j}R_{ext}(j)}{\max_{j}R_{ext}(j)-\min_{j}R_{ext}(j)}$ (9)
@@
    $R_{agg}(i)=\frac{\tilde{R}_{ELHSR}(i)+\tilde{R}_{ext}(i)}{2}$ (10)
@@
    $i^{*}=\arg\max_{i\in\{1,...,N\}}R_{agg}(i)$ (11)
These combination strategies allow us to leverage both internal (ELHSR) and external reward signals.
Appendix C Implementation Details
@@
C.1 Hardware and Software
We conduct our experiment in a system with 8x NVIDIA A100 GPUs (80GB each) and an Intel(R) Xeon(R) Gold 6346 CPU @ 3.10GHz with 1.0TB of CPU memory (large CPU memory is unnecessary). We used the PyTorch framework [Paszke et al., 2019] for implementing our models and training procedures.
C.2 Details of the Preliminary Experiments
@@
  1. 1.
    Hidden State Extraction: We use the Llama-3.1-8B-Instruct model [Grattafiori et al., 2024] and extract the hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ during the generation of reasoning paths on the MATH dataset [Hendrycks et al., 2021].
  2. 2.
@@
  3. 3.
    Dimensionality Reduction (PCA): We apply Principal Component Analysis (PCA) [Wold et al., 1987] to reduce the dimensionality of $\bar{h}^{\ell}$ to 50 dimensions. This step helps to mitigate overfitting that we observed with direct application of LDA.
  4. 4.
    Classification (LDA): We train a Linear Discriminant Analysis (LDA) classifier [Balakrishnama and Ganapathiraju, 1998] on the reduced representations to predict whether the reasoning path is correct or incorrect.
The entire pipeline is linear to validate that hidden states of the reasoning paths encode a linear representation. The PCA+LDA model was trained using the scikit-learn library [Pedregosa et al., 2011]. We adopted the default settings in scikit-learn.
C.3 ELHSR Training Details
The ELHSR model was trained using the AdamW optimizer [Loshchilov and Hutter, 2019] with the following hyperparameters:

### v3

@@ 5.8. Logit-Based Training for Closed-Source LLMs
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
@@
While SWIFT offers significant advantages in terms of efficiency and performance, making it well-suited for deployment by large model providers and end-users, a limitation arises when working with closed-source LLMs where hidden states are inaccessible. To overcome this constraint, we leverage the fact that some closed-source models still provide access to their logits, such as GPT-3.5-turbo (OpenAI, 2023a) and GPT-4 (Achiam et al., 2023), as demonstrated by Finlayson et al. (2024) and OpenAI (2023b). We train SWIFT directly on these logits, resulting in a model with a parameter count that is orders of magnitude smaller than traditional reward models - and even smaller than SWIFT trained on all layers. This enables seamless deployment on personal devices such as smartphones and personal computers. As demonstrated in Table 6 and Appendix Table 12, the logit-based SWIFT maintains strong performance. Furthermore, as shown in Table 1, its performance even surpasses that of many baselines. These results highlight the versatility and adaptability of SWIFT, making it a viable solution even when access to hidden states is restricted.
Concretely, with BoN@64, logit-only SWIFT reaches 50.6/55.8/57.4 on MATH, 81.0/91.2/90.2 on GSM8K, and 65.2/78.0/74.0 on AQuA_RAT for Llama-3.2-3B/Llama-3.1-8B/Ministral-8B, respectively, demonstrating that informative correctness signals can be extracted even from logit access alone.
5.9. Combining SWIFT with External Reward Models
@@
Figure 5. Average performance of combined SWIFT and external reward models across different datasets and task-performing models, for both rank selection and scaled averaging. See Appendix Figure 8 and Figure 9 for full details.
Average performance of combined SWIFT and external reward models across different datasets and task-performing models.
@@
SWIFT offers a highly parameter-efficient approach to reward modeling, enabling its seamless integration with existing external reward models. Given SWIFT’s minimal parameter footprint (less than 0.005% of baseline models), the additional computational overhead introduced by incorporating SWIFT is negligible compared to relying solely on the external reward model. We explore two straightforward methods for combining SWIFT with external rewards: rank selection and scaled averaging. In rank selection, we choose the sample with the best rank as determined by either SWIFT or the external reward model. In scaled averaging, we compute the average of the rewards from both models after normalizing them to the [0, 1] range. Detailed descriptions of these methods are provided in Appendix C.
@@
We evaluated the effectiveness of these combination strategies by integrating SWIFT with various traditional reward models across different datasets and task-performing models, assessing performance using best-of-N sampling. Figure 5 presents the average results across different datasets for both ranking and scaling methods, applied to various task-performing models. Detailed results for each dataset and task-performing model can be found in Appendix Figure 8 and Figure 9. Our findings indicate that combining SWIFT with external reward models can lead to further improvements in accuracy, particularly on the GSM8K and AQuA-RAT datasets. The MATH dataset did not exhibit significant performance gains, potentially because SWIFT already significantly outperforms the external reward models on this benchmark. This performance boost is likely due to SWIFT and external reward models capturing different types of errors, resulting in improved overall accuracy-particularly with a higher number of best-of-N reasoning paths. This combination strategy offers a pathway to further performance gains without incurring significant computational overhead.
5.10. Analysis on SWIFT
@@
To investigate the mechanism of SWIFT, we present illustrative examples of its scoring of reasoning paths in Appendix F. The examples with the highest and lowest gating $\times$ reward values are highlighted, along with their corresponding numerical scores. Notably, SWIFT assigns relatively high or low gating $\times$ reward values to specific numerical and mathematical symbols, as well as particular tokens such as “answer”, “boxed” and the special token ‘<|eot_id|>’. This behavior suggests that SWIFT possesses the ability to identify and weigh key components within the reasoning process.
In addition, we also conduct ablation study on the gating mechanism of SWIFT. SWIFT includes a lightweight token-level gating module that re-weights token representations before aggregating them into a path-level score. We ablate this component by removing gating while keeping the same training budget and evaluation protocol. On MATH, gating yields consistent gains across all BoN settings. For BoN@64, accuracies improve from 49.4 to 53.6 (Llama-3.2-3B), 57.6 to 62.6 (Llama-3.1-8B), and 61.4 to 62.8 (Ministral-8B). Similar trends hold at BoN@4 and BoN@16, confirming that gating is beneficial across varying numbers of candidate paths. This token-level selection mechanism allows SWIFT to emphasize semantically decisive portions of reasoning traces, such as intermediate steps, numerical computations, and conclusion markers, while downweighting boilerplate text. The details can be found in Table 8.
To better understand where SWIFT’s gains come from, we further break down improvements by MATH sub-domain (Ministral-8B as the task LLM; BoN@64). SWIFT improves over Math-Shepherd-7B across all sub-domains, with gains ranging from +2.44 points (Geometry) to +10.52 points (Counting & Probability), and consistently strong gains on Algebra (+8.87), Intermediate Algebra (+10.31), and Number Theory (+9.68). See Table 9 for the full breakdown.
We also compare alternative objectives for training SWIFT on MATH, including hinge-style and preference-style losses. Across three task LLMs (Llama3-3B/Llama3-8B/Ministral-8B) and three BoN settings (@4, @16, @64), cross-entropy achieves the best average accuracy (57.5), followed by DPO (57.0), hinge loss (56.8), InfoNCA (55.1) and NCA (54.4). The relatively small gap between the top three suggests that SWIFT is robust to the choice of training signal. Cross-entropy remains our default due to its consistent edge and simpler implementation. The detailed experiment is illustrated in Appendix E.1.
6. Conclusion
@@
In this paper, we presented SWIFT, a novel reward modeling approach for LLMs that leverages the rich information contained in their internal hidden states. SWIFT offers both high efficiency and effectiveness, and it scales well during both training and inference. We further showed that comparable performance can be achieved by training SWIFT on a subset of layers or even on logits alone, significantly reducing computational costs and making it more versatile. The successful integration of SWIFT with standard reward models underscores its flexibility and potential.
Admittedly, this work has some limitations. SWIFT’s reliance on access to the LLM’s hidden states or logits may reduce its applicability to commercial LLMs that do not expose this information to users. However, its efficiency and ease of deployment make SWIFT especially well-suited to LLM providers themselves. In addition, SWIFT must be trained separately for each LLM as hidden-state dimensionality and layer count vary across models. Nonetheless, given its small model size and minimal data requirements, the associated computational overhead is negligible, especially when compared to the cost of training the task-performing LLM itself.
Acknowledgements.
@@
This work was supported by the National Science Foundation (NSF) under Grant No. III-2106758. We gratefully acknowledge Mrs. Jiahe Jin from Zhiyuan College, Shanghai Jiao Tong University, for her invaluable assistance in creating the initial version of our illustrations. We also thank Mr. Zhihui Xie from The University of Hong Kong for offering precious suggestions.
@@
Appendix A Relevance to KDD Community
Our work is highly relevant to the KDD community and aligns directly with the topic of Modern AI and Big Data of the KDD 2026 Research Track, since according to the description of the call for paper page in the official website, deep representation learning, generation and reasoning are all contained in this topic. We address a critical challenge in the large-scale deployment of Large Language Models (LLMs) by shifting the focus of the data source for reward modeling. Instead of relying on computationally prohibitive, text-based reward models, our work introduces SWIFT, a technique that mines an LLM’s internal hidden states to extract an intrinsic signal of reasoning correctness. This approach embodies the principles of deep representation learning by leveraging the rich information encoded within a model’s internal layers. Furthermore, it directly contributes to advancements in generation and reasoning by offering a lightweight and highly efficient method to enhance output quality. By demonstrating that valuable, actionable knowledge can be discovered from a model’s internal states, which is a novel data source, our paper presents a significant contribution to the efficient and scalable application of large language models (LLMs), a core topic of interest to the KDD community.
Appendix B Baseline Reward Models Description
@@
  • EurusRM-7B: EurusRM-7B is a 7B parameter reward model trained on a mixture of UltraInteract (Yuan et al., 2025), UltraFeedback (Cui et al., 2024), and UltraSafety datasets (Guo et al., 2024), with a focus on improving reasoning performance. EurusRM-7B stands out as the best 7B RM overall and achieves similar or better performance than much larger baselines. Particularly, it outperforms GPT-4 in certain tasks.
  • Skywork-Reward-Llama-3.1-8B-v0.2: This 8B parameter reward model is built on the Llama-3 architecture (Grattafiori et al., 2024) and trained on a curated dataset of 80K high-quality preference pairs. It excels at handling preferences in complex scenarios, including challenging preference pairs, and span various domains such as mathematics, coding, and safety. As of October 2024, Skywork-Reward-Llama-3.1-8B-v0.2 ranks first among 8B models.
  • Starling-RM-7B: Starling-RM-7B is a reward model trained from Llama2-7B-Chat (Touvron et al., 2023), following the method of training reward models in the instructGPT paper (Ouyang et al., 2022). It is trained with the Nectar preference dataset (Zhu et al., 2024), which is based on GPT-4’s preferences.
  • UltraRM-13B: UltraRM is a reward model initialized by Llama2-13B (Touvron et al., 2023) and fine-tuned on the UltraFeedback dataset (Cui et al., 2024). It achieves state-of-the-art performance among open-source reward models on public preference test sets.
  • RLHFlowRM-8B-Deepseek: RLHFlowRM-8B-Deepseek is trained from Llama-3.1-8B-Instruct on Deepseek generated data. It is evaluated on GSM8K and MATH dataset and demonstrated significant performance gains.
  • Math-Shepherd-7B: Math-Shepherd is trained with automatic process annotation. It is benchmarked on GSM8K and MATH dataset and substantially boosted the task-performing LLM’s performance.
Appendix C Detailed Approach for Combining SWIFT with External Reward Models
In the Best-of-N (BoN) setting, we explore two strategies for combining the SWIFT reward ($R_{SWIFT}$) with an external reward signal ($R_{ext}$), typically from a traditional reward model. We have $N$ reasoning paths.
  1. (1)
    Ranking-based Combination: Let $rank_{SWIFT}(i)$ and
    $rank_{ext}(i)$ be the ranks of the $i$-th reasoning path according to SWIFT and the external reward model, respectively (lower rank is better). The relative rank is:
    (6) $rank_{rel}(i)=\max(rank_{SWIFT}(i),rank_{ext}(i))$
@@ Appendix C Detailed Approach for Combining SWIFT with External Reward Models
    (8) $\tilde{R}_{SWIFT}(i)=\frac{R_{SWIFT}(i)-\min_{j}R_{SWIFT}(j)}{\max_{j}R_{SWIFT}(j)-\min_{j}R_{SWIFT}(j)}$
    (9) $\tilde{R}_{ext}(i)=\frac{R_{ext}(i)-\min_{j}R_{ext}(j)}{\max_{j}R_{ext}(j)-\min_{j}R_{ext}(j)}$
@@
    (10) $R_{agg}(i)=\frac{\tilde{R}_{SWIFT}(i)+\tilde{R}_{ext}(i)}{2}$
@@
    (11) $i^{*}=\arg\max_{i\in\{1,...,N\}}R_{agg}(i)$
These combination strategies allow us to leverage both internal (SWIFT) and external reward signals.
Appendix D Implementation Details
@@
D.1. Hardware and Software
We conduct our experiment in a system with 8x NVIDIA A100 GPUs (80GB each) and an Intel(R) Xeon(R) Gold 6346 CPU @ 3.10GHz with 1.0TB of CPU memory (large CPU memory is unnecessary). We used the PyTorch framework (Paszke et al., 2019) for implementing our models and training procedures.
D.2. Details of the Preliminary Experiments
@@
  1. (1)
    Hidden State Extraction: We use the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) and extract the hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ during the generation of reasoning paths on the MATH dataset (Hendrycks et al., 2021).
  2. (2)
@@
  3. (3)
    Dimensionality Reduction (PCA): We apply Principal Component Analysis (PCA) (Wold et al., 1987) to reduce the dimensionality of $\bar{h}^{\ell}$ to 50 dimensions. This step helps to mitigate overfitting that we observed with direct application of LDA.
  4. (4)
    Classification (LDA): We train a Linear Discriminant Analysis (LDA) classifier (Balakrishnama and Ganapathiraju, 1998) on the reduced representations to predict whether the reasoning path is correct or incorrect.
The entire pipeline is linear to validate that hidden states of the reasoning paths encode a linear representation. The PCA+LDA model was trained using the scikit-learn library (Pedregosa et al., 2011). We adopted the default settings in scikit-learn.
D.3. SWIFT Training Details
The SWIFT model was trained using the AdamW optimizer (Loshchilov and Hutter, 2019) with the following hyperparameters: